OpenStudy (beth12345):

if h(x)=f(x)g(x) and f(x)=2x+5, determine g(x) a) h(x)=3x-1/x+7 b h(x)=3x-1/the square root of x+6

OpenStudy (beth12345):

a)\[h(x) = \frac{ 3x-1 }{ x+7 }\]

OpenStudy (beth12345):

and b) \[h(x)=\frac{ 3x-1 }{ \sqrt{x+6} }\]

OpenStudy (anonymous):

You can start by writing after the equation in part (a) for h(x) that it also equals f(x)g(x) which equals (2x+5)g(x) according to the given

OpenStudy (anonymous):

\[h(x) = \frac{ 3x-1 }{ x+7 } = (2x+5)g(z)\]

OpenStudy (anonymous):

Then divide both sides by (2x+5) to isolate g(x) (sorry, I accidentally put g(z) in the last equation)

OpenStudy (anonymous):

Then simplify the fraction expression if you can

OpenStudy (anonymous):

Let me know if this makes sense... :)

OpenStudy (beth12345):

so would it be like \[g(x)=\frac{( \frac{ 3x-1 }{ x+7 }) }{ 2x+5 }\] and then simplify it?

OpenStudy (anonymous):

yes :) You can start simplifying by "moving" that (x+7) term down to the bottom, like: g(x) = (3x-1) / [ (x+7)(2x+5) ]

OpenStudy (anonymous):

But that's the idea... they give you h(x) = f(x)g(x) and they give you f(x) and h(x) as expressions, so you just substitute those expression in where they belong, and you are left with g(x) and a bunch of x terms and constants to simplify to obtain g(x).

OpenStudy (anonymous):

In basic algebra, you might substitute in values for variables to solve for one variable. Same idea here, but you sub in function/expressions for the known functions to solve for the function you don't know, in this case, g(x)

OpenStudy (beth12345):

ok

OpenStudy (beth12345):

i cant really get how to simplify it past \[g(x)=(3x-1)/3^{2}+19x+35\]

OpenStudy (anonymous):

where you have 3^2 on the bottom, it's actually x^2, but yeah, I see what you mean... I'm not seeing a simplification yet.

OpenStudy (anonymous):

That may be ok... nothing says g(x) must be a simple expression. Although usually if things get too messy, it's good to look for an simple sign error or something.

OpenStudy (beth12345):

k, so would i do the second one the same way to?

OpenStudy (anonymous):

Sorry, stepped away... yes, the second one is the same exact idea, and it too looks messy with likely no way to simplify that square root term... don't worry that the g(x) expressions don't "look nice", just simplify what you can. They gave you messy h(x) expressions... that's what's causing it.

OpenStudy (beth12345):

ok, there was another one to, it had all of the same h(x) things and it was h(x)=6x+15

OpenStudy (beth12345):

so then i got 6x+15=(2x+5)(g(x))

OpenStudy (anonymous):

Remember, h(x) = f(x)g(x) according to the given. So, 6x+15 = (2x+5)g(x). Solve for g(x) :) You beat me to it :)

OpenStudy (anonymous):

Note that you finally (!!) can factor the expression to simplify it...

OpenStudy (beth12345):

and then \[\frac{ 6x+15 }{ 2x+5 } = g(x)\]

OpenStudy (anonymous):

Always look at the denominator to see if it might be an easy factor of the numerator

OpenStudy (anonymous):

factor out 3 on top --> 3(2x +5) and cancel with the denominator term leaving g(x) = 3

OpenStudy (anonymous):

Then say "duh" because when you look at h(x), it's just 3 times f(x), so g(x) = 3

OpenStudy (anonymous):

I didn't see that at first

OpenStudy (beth12345):

haha thanks, ya i was like ooooh

OpenStudy (anonymous):

Nice!

OpenStudy (beth12345):

ok, i get how to do this now, thanks allot :D

OpenStudy (anonymous):

Glad to help!!

OpenStudy (beth12345):

glad for the help :P lol